OpenStudy (nbarrera):

I Need Heeelp+Medal Which number is between 1/6 and 0.25 A.1/7 B.3/16 C.7/24 D.2/7

OpenStudy (nbarrera):

I need help here

OpenStudy (kiakat):

Change the 1/6 into a decimal

OpenStudy (nbarrera):

anyone?

OpenStudy (kiakat):

Either that or the .25 into a fraction, either will work

OpenStudy (kiakat):

actually for this specific question it'll be easier to change all of your fractions into decimals (even the ones for your answers)

OpenStudy (anonymous):

1/6 is .16666.. so you are looking for a number between .1666... and .25.

OpenStudy (nbarrera):

is 1/6 =0.16666667

OpenStudy (anonymous):

what do you think the answer is?

OpenStudy (nbarrera):

also need to convert 0.25 in fraction

OpenStudy (anonymous):

1/4

OpenStudy (nbarrera):

so 1/4 =0.25 in decimal

OpenStudy (nbarrera):

So whats between 1/6 and 1/4?

OpenStudy (nbarrera):

A.1/7 B.3/16 C.7/24 D.2/7

OpenStudy (javalos01):

I got this dont worry

OpenStudy (anonymous):

5 / 24

OpenStudy (anonymous):

((1/4) + (1/6)) / 2 ((3/12) + (2/12)) / 2 (5/12) / 2 5 / 24

OpenStudy (nbarrera):

But in the answer choice i dont have 5/24

OpenStudy (anonymous):

yeah i just saw that. Hold on let me try this again. Let me redo the work.

OpenStudy (javalos01):

iIl try too

OpenStudy (nbarrera):

The lesson im in is about Terminating and Repeating Numbers

OpenStudy (javalos01):

c

OpenStudy (javalos01):

In my opinion I think its C. :D

OpenStudy (anonymous):

Got to find a number that is 0.166667 & 0.25 = 1/6 = 1 ÷ 6 = 0.166667 A) 1/7 = 1/7 = 1 ÷ 7 = 0.142857 B) 3/16 = 3/16 = 3 ÷ 16 = 0.1875 c) 7/24 = 7/24 = 7 ÷ 24 = 0.291667 D) 2/7 = 2/7 = 2 ÷ 7 = 0.285714 Its b.

OpenStudy (nbarrera):

Ok Thanks hope it's right.

OpenStudy (anonymous):

look at it this way.. 3/16 is 0.18... The dots means it goes on. Lets just take the two numbers that come after the dec of all them. 1/7 is 0.14.. 3/16 is 0.18.. 2/7 is0.28.. 7/24 is 0.29.. Put them in order and pick what one should go in between 0.16.. & 0.25 it would be 0.18..

OpenStudy (nbarrera):

Yes and Can you check if my answer is correct? Which number is between 0.5 and 5/8? A.8/28 B.1/4 C.19/28 D.15/28<==my answer

OpenStudy (anonymous):

yeah post it and i will. tag me in it tho.

OpenStudy (anonymous):

Did you post it?

OpenStudy (nbarrera):

not yet wait

OpenStudy (javalos01):

was it c